OpenStudy (anonymous):

Which of these is the area of a sector of a circle with r = 18”, given that its arc length is 6π? A. 54.00 in^2 B. 113.10 in^2 C. 169.65in^2 D. 339. 29 in^2

OpenStudy (anonymous):

\[\frac{\text{arc length}}{\text{sector area}}=\frac{\text{total circumference}}{\text{total area}}~~\implies~~\frac{6\pi}{\text{sector area}}=\frac{2\pi r}{\pi r^2}\] Given that \(r=18\), you can easily compute the sector area.